# Learning-based Distributed Model Predictive Control using Multi-Agent Bayesian Optimization

### v1

@@ Abstract
This paper presents a fusion of Multi-agent Bayesian Optimization (MABO) and Distributed Model Predictive Control (DMPC) aiming at learning the DMPC schemes with imperfect local models in a distributed manner. In the proposed method, we use a dual-decomposition method for a DMPC and leverage an Alternating Direction Method of Multipliers (ADMM)-based MABO for coordinated learning of the parameterized DMPC scheme to improve the closed-loop performance of the local MPC schemes even if their models cannot capture the real multi-agent system perfectly.
@@ I Introduction
Model Predictive Control (MPC) is a control strategy that uses a predictive model to compute optimal control actions by solving a constrained optimization problem over a receding horizon [1]. It is widely used in applications requiring constraint satisfaction and multi-variable control, including autonomous vehicles, chemical processes, and energy systems. However, due to uncertainties and unknown dynamics, accurate models of dynamical systems are often difficult to obtain. Even if accurate models are available, they may be in general too complex to be used in MPC schemes. Consequently, the performance degradation often occurs due to inaccurate models used in MPC. Furthermore, choosing a suitable open-loop cost function and constraints to achieve the desired closed-loop performance while guaranteeing safety remains challenging.
@@
A common approach to managing large-scale nonlinear networked systems is to design local controllers that neglect interactions between subsystems. However, this often results in significantly degraded overall system performance. While centralized control could achieve better global performance, it is typically impractical due to unavoidable communication limitations, the complexity of nonlinear systems, and the high number of decision variables [2]. To address these challenges, extensive research has focused on structured control systems, including decentralized and distributed architectures. However, the primary challenge in a decentralized architecture lies in the lack of communication between controllers. This limitation can lead to degraded closed-loop performance and, in some instances, even result in instability. In contrast, a distributed control system allows communication between controllers, enabling better coordination and potentially improved performance [3].
In the context of distributed control systems, Distributed MPC (DMPC) is a well-known control approach to dealing with interconnected systems. DMPC has desirable features for distributed control, such as handling multi-variable interactions, constraints on control inputs, and system states. Dual-decomposition and Alternating Direction Method of Multipliers (ADMM) are two efficient methods for solving DMPC problems where a coupled constraint between the agents can be formulated as a dual problem [4, 5]. However, the inaccurate local MPC models significantly affect the performance of DMPC. It can also be challenging to chose parameters for the cost and constraint functions for each agent.
@@
Machine learning has emerged as a promising approach to address this issue by enabling data-driven model learning for MPC [6, 7]. However, the closed-loop control performance is not directly related to the model fitting so that the control objectives may not be satisfied even if the learned model can capture the real plant accurately. To address this issue, the methods in [8, 9, 10, 11] established the concepts behind the fusion of Markov Decision Process (MDP) and MPC so that an MPC scheme can deliver the same optimal policy as MDP by modifying the terminal and stage cost functions of MPC. Then, several Reinforcement Learning (RL) methods were developed based on the MPC schemes to learn the corresponding MPC cost functions aiming at improving the closed-loop performance. More precisely, parametrizing the cost function offers an alternative approach to model learning, motivated by the inherent connection between model-based predictions and the MPC cost function.
The integration of Bayesian Optimization (BO) with MPC has recently represented a powerful paradigm for addressing complex control problems in various domains. By enabling adaptive, data-driven, and efficient control strategies, this combination holds the potential for widespread application in industries ranging from robotics to energy and beyond [12, 13, 14]. BO is a probabilistic optimization approach designed for black-box functions that are expensive to evaluate. The methodology revolves around the construction of a surrogate model, often a Gaussian Process (GP), to approximate the target function [15].
@@
Combining BO with MPC leverages the strengths of both methodologies: BO’s ability to optimize black-box functions and MPC’s predictive control capabilities. MPC performance hinges on selecting parameters like prediction horizon, weighting matrices, and constraints. BO then enables systematic tuning by considering the closed-loop system’s performance as a black-box objective function [16, 17]. The BO’s surrogate modeling also can be used to capture and update system uncertainties in real-time, enabling adaptive MPC strategies for nonlinear or time-varying systems [18]. Authors in [19] proposed a safe and stability-informed BO for MPC cost function learning in which a parameterized MPC scheme is safely adjusted to achieve its best closed-loop performance in the presence of a model-mismatch between the MPC model and real plant. A high-dimensional BO framework for sample-efficient MPC tuning was proposed in [20].
@@
However, in the context of multi-agent systems, the combined BO-MPC control approaches above do not account for the coupling and the interactions between the different agents, i.e., through a DMPC scheme. In this paper, we then propose to learn a DMPC scheme using a Multi-Agent BO (MABO) framework aiming at improving the best closed-loop performance for each local MPC scheme in the presence of model-mismatch. We show the fusion of DMPC and multi-agent Markov Decision Processes (MDPs) so that the local MPC schemes obtained from a dual-decomposition method can capture the local value functions associated with the multi-agent MDP. In the proposed learning-based DMPC, we first show how the local cost functions associated with the local MPC schemes can be modified such that the local optimal policies are delivered. We then propose to practically learn a parameterized DMPC by a coordinated learning mechanism based on the ADMM-based MABO.
@@
This paper is organized as follows. In Section II, a distributed MPC scheme based on a dynamic dual-decomposition method is described. The fusion of DMPC and multi-agent MDPs is detailed in III. In Section IV, we describe the coordinated-based BO algorithm for learning the DMPC schemes. A numerical example is then illustrated in Section V to show the performance of the proposed MABO-DMPC and finally we conclude the paper in Section VI.
II Distributed MPC
In this section, we provide formulations for the centralized control problem and the use of a dual-decomposition method to break the centralized problem into several local optimization problems.
II-A Networked Optimal Control System
@@
This work considers a distributed networked optimal control system comprising a set of $\mathcal{M}$ interconnected dynamical subsystems, each denoted as: =$\mathcal{M}=\{\Sigma_{1},\dots,\Sigma_{|\mathcal{M}|}\}$, where $m:=|\mathcal{M}|$ represents the cardinality of the set $\mathcal{M}$. For each subsystem $\Sigma_{i}$, we define subsystem $\Sigma_{j}$ as a neighbor of $\Sigma_{i}$ if the two subsystems are subject to coupled constraints and/or a coupled cost function. Let us consider a Distributed Control System (DCS), which is defined as a graph $\mathcal{G}(\mathcal{M},\mathcal{E}_{c})$, consisting of a set of nodes $\mathcal{M}$ corresponding to the subsystems, and edges $\mathcal{E}_{c}$, representing the interconnections between the subsystems through a coupling cost function. The set $\mathcal{M}_{i}:=\{\Sigma_{j}\,|\,(i,j)\in\mathcal{E}_{c}\text{ or }(j,i)\in\mathcal{E}_{c},\,i\neq j\}$ then represents the set of subsystems $\Sigma_{j}$ that are interconnected with $\Sigma_{i}$.
II-B Centralized Control
@@
Let the state and input of the agent $i$ be denoted by $\boldsymbol{\mathrm{x}}_{i}$ and $\boldsymbol{\mathrm{u}}_{i}$, respectively. We consider a deterministic model of each agent as $\boldsymbol{\mathrm{x}}_{i}^{k+1}=\boldsymbol{\mathrm{f}}_{i}(\boldsymbol{\mathrm{x}}_{i}^{k},\,\boldsymbol{\mathrm{u}}_{i}^{k})$, where $\boldsymbol{\mathrm{x}}_{i}^{k}\in\mathbb{R}^{n_{x_{i}}},\boldsymbol{\mathrm{u}}_{i}^{k}\in\mathbb{R}^{n_{u_{i}}}$. Let $\boldsymbol{\mathrm{x}}^{k}=\left[\boldsymbol{\mathrm{x}}_{1}^{k},\ldots,\boldsymbol{\mathrm{x}}_{m}^{k}\right]^{\top}$ and $\boldsymbol{\mathrm{u}}^{k}=\left[\boldsymbol{\mathrm{u}}_{1}^{k},\cdots,\boldsymbol{\mathrm{u}}_{m}^{k}\right]^{\top}$ be the augmented state and control input vectors of multi-agent system, respectively. The corresponding dimensions then read as $\boldsymbol{\mathrm{x}}^{k}\in\mathbb{R}^{n_{x}},\boldsymbol{\mathrm{u}}^{k}\in\mathbb{R}^{n_{u}}$, $n_{x}=\sum_{i=1}^{m}n_{x_{i}},n_{u}=\sum_{i=1}^{m}n_{u_{i}}$. A networked control scheme can be based on a centralized optimization problem or set of local problems that need to be solved at time instant $k$. The centralized optimization problem is then formulated as
$\min_{\hat{\boldsymbol{\mathrm{x}}},\hat{\boldsymbol{\mathrm{u}}}}\sum_{i=1}^{m}T_{i}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{k+N},\hat{\boldsymbol{\mathrm{w}}}_{i}^{k+N}\right)+\sum_{\ell=k}^{k+N-1}l_{i}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{w}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{u}}}_{i}^{\ell}\right)$ (1a)
@@
where $W_{ij}\left(\hat{\boldsymbol{\mathrm{x}}}_{j}^{\ell}\right)_{(j,i)\in\mathcal{E}_{c}}\in\mathbb{R}^{\Sigma_{(j,i)\in\mathcal{E}_{c}}n_{j}}$ denotes the tuple of the state vector of all the subsystems that can influence subsystem $i$. $N$ is the prediction horizon and $T_{i}$, $l_{i}$, $h_{i}$ and $g_{i}$ denote the respective terminal cost, stage cost, mixed inequality constraint, and input inequality constraint for agent $i$.
@@
In this paper, we assume that the subsystems of a multi-agent system have the same dimension of the state-space while their dynamics can be different yielding a heterogeneous multi-agent system.
@@
In this paper, we consider a multi-agent system with fully coupled subsystems such that each agent $\Sigma_{i}$ is affected by all the other agents $\mathcal{M}_{i}$. According to Assumption 1, we then have that $\hat{\boldsymbol{\mathrm{w}}}_{i}^{\ell}\in\mathbb{R}^{(m-1)n_{x_{i}}}$ in (1d).
Solving (1) yields a sequence of optimal input predictions and corresponding state predictions as
@@
II-C Dynamic Dual-Decomposition
Taking into account the coupling constraints (1d) of the centralized optimization problem (1), one can introduce the slack variable $\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}$ capturing the effect of other agents on the agent $i$ through $W_{ij}\left(\hat{\boldsymbol{\mathrm{x}}}_{j}^{k+\ell}\right)_{(j,i)\in\mathcal{E}_{c}}$. The centralized optimization scheme can then be rewritten as
@@ II-C Dynamic Dual-Decomposition
$+\left(\boldsymbol{\mathrm{\mu}}_{i}^{\ell}\right)^{\top}\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}-\sum_{j=1,j\neq i}^{m}\left(\boldsymbol{\mathrm{\mu}}_{ji}^{\ell}\right)^{\top}W_{ji}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell}\right).$
where $\boldsymbol{\mathrm{\mu}}_{i}^{\ell}=\left(\boldsymbol{\mathrm{\mu}}_{ij}^{\ell}\right)_{(j,i)\in\mathcal{E}_{c}}\in\mathbb{R}^{\Sigma_{(j,i)\in\mathcal{E}_{c}}n_{j}}$. The local MPC scheme is then formulated as
@@
$\qquad=\operatorname*{arg\,min}_{\hat{\boldsymbol{\mathrm{x}}}_{i},\hat{\boldsymbol{\mathrm{u}}}_{i},\bar{\boldsymbol{\mathrm{w}}}_{i}}T_{i}^{\boldsymbol{\mathrm{\mu}}^{(I)}}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{k+N},\bar{\boldsymbol{\mathrm{w}}}_{i}^{k+N}\right)$ (5a)
$\qquad\qquad\qquad\qquad\qquad+\sum_{\ell=k}^{k+N-1}L_{i}^{\boldsymbol{\mathrm{\mu}}^{(I)}}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{u}}}_{i}^{\ell},\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}\right)$
@@
where $\beta_{i}>0$ is the stepsize. The stopping criteria then holds ($\boldsymbol{\mathrm{\mu}}_{i}^{I}\rightarrow\boldsymbol{\mathrm{\mu}}_{i}^{\star}$) if either the number of iterations exceeds $I\geq I_{\text{max}}$ or one of the following conditions is satisfied.
@@
for some positive threshold $\epsilon_{1},\epsilon_{2}$. It is worth noting that each of the two conditions above can be used as a valid stopping criterion if local MPC models can accurately capture the real subsystems. Consequently, we denote by $\hat{\boldsymbol{\mathrm{\mu}}}_{i}^{\star}$ an imperfect convergence of $\boldsymbol{\mathrm{\mu}}_{i}^{\star}$ when dealing with imperfect MPC models.

### v2

# Distributed Model Predictive Control Design for Multi-agent Systems via Bayesian Optimization

@@ Abstract
This paper introduces a new approach that leverages Multi-agent Bayesian Optimization (MABO) to design Distributed Model Predictive Control (DMPC) schemes for multi-agent systems. The primary objective is to learn optimal DMPC schemes even when local model predictive controllers rely on imperfect local models. The proposed method invokes a dual decomposition-based distributed optimization framework, incorporating an Alternating Direction Method of Multipliers (ADMM)-based MABO algorithm to enable coordinated learning of parameterized DMPC schemes. This enhances the closed-loop performance of local controllers, despite discrepancies between their models and the actual multi-agent system dynamics. In addition to the newly proposed algorithms, this work also provides rigorous proofs establishing the optimality and convergence of the underlying learning method. Finally, numerical examples are given to demonstrate the efficacy of the proposed MABO-based learning approach.
@@ I Introduction
Networked systems have grown significantly in scale and complexity, leading to increasingly demanding requirements for model-based control accuracy and real-time performance. For such large-scale network systems, complex optimization algorithms are commonly employed to perform desired tasks while minimizing a priori defined objective functions and satisfying operational constraints [1]. Optimization-based control of networked systems then remains a challenging problem when a complex model of multi-agent system is exploited within the optimization framework. Although simplified models can reduce the complexity of the model-based optimization algorithms, the inaccuracies can significantly degrade the performance of multi-agent control system.
@@
A common approach to managing large-scale networked systems is to design local controllers that neglect interactions between the subsystems involved. However, this often results in an overall degradation in the system performance. Although centralized control could achieve better global performance, it is generally impractical due to limitations in communication, the complexity of nonlinear systems, and the large number of decision variables [2]. To address these challenges, extensive research has focused on structured control systems, including decentralized and distributed architectures. However, the primary challenge in a decentralized architecture lies in the lack of communication between controllers, which can lead to degraded closed-loop performance and, in some instances, even result in instability. In contrast, a distributed control system allows communication between controllers, enabling better coordination and potentially improved performance [3].
@@
In the context of distributed control systems, distributed MPC (DMPC) is a well-researched optimal control approach that can handle interconnected systems and multi-variable interactions. Dual decomposition and Alternating Direction Method of Multipliers (ADMM) are two efficient methods for solving DMPC problems, where a coupled constraint between agents can be formulated as a dual problem [4, 5]. However, inaccurate local MPC models adversely affect the performance of DMPC. It can also be challenging to choose parameters for the cost and constraint functions for each agent. Machine learning-based approaches have recently emerged to address the aforementioned issue by enabling data-driven model learning for MPC [6, 7]. However, closed-loop control performance is not directly related to the model fitting, so that the control objectives may not be satisfied even if the learned model can accurately capture the real plant. To address this issue, the methods proposed in [8, 9, 10, 11, 12] established tools to fuse Markov Decision Process (MDP) and MPC so that an MPC scheme can deliver the same optimal policy as MDP by modifying the terminal and stage cost functions of the MPC. Then, several Reinforcement Learning (RL)-based methods were developed to learn the corresponding MPC cost functions aimed at improving closed-loop performance. It was shown that parameterizing the cost function can offer an alternative approach to model learning, motivated by the inherent connection between model-based predictions and the MPC cost function.
@@
The integration of Bayesian Optimization (BO) with MPC has recently introduced a powerful paradigm for addressing complex control problems in various domains. By enabling adaptive, data-driven, and efficient control strategies, this combination holds the potential for many applications ranging from robotics to energy systems and beyond [13, 14, 15]. BO is a probabilistic optimization approach designed for black-box functions that are expensive to evaluate. The methodology revolves around the construction of a surrogate model, often a Gaussian Process (GP), to approximate the target function [16]. Appropriate fusion of BO with MPC leverages the strengths of both; while BO has the ability to optimize black-box functions, MPC offers predictive control capabilities. MPC performance hinges on the proper selection of parameters such as prediction horizon, weighting matrices, and constraints. BO then enables systematic tuning by considering the closed-loop system performance as a black-box objective function [17, 18]. The BO’s surrogate modeling can also be used to capture and update system uncertainties in real time, enabling adaptive MPC strategies for nonlinear or time-varying systems [19]. Authors in [20] proposed a safe and stability-informed BO for MPC cost function learning, in which a parameterized MPC scheme is safely adjusted to deliver the best closed-loop performance in the presence of a model mismatch between the MPC model and the real plant. A high-dimensional BO framework for sample-efficient MPC tuning was also proposed recently in [21].
@@
Despite recent efforts in utilizing BO for control design purposes, in the context of multi-agent systems, the combined BO-MPC control approaches above do not account for the coupling and the interactions between different agents, i.e., through a DMPC scheme. To address that, in this paper, we propose to learn a DMPC scheme using a Multi-Agent BO (MABO) framework aiming at improving the optimal closed-loop performance for each local MPC scheme in the presence of model mismatch. We formulate the fusion of DMPC and multi-agent Markov Decision Processes (MDPs) so that the local MPC schemes obtained from a dual-decomposition method can capture the local value functions associated with the multi-agent MDP. In the proposed learning-based DMPC, we first show how the local cost functions associated with the local MPC schemes can be modified to deliver local optimal policies. We then propose to practically learn a parameterized DMPC by a coordinated learning mechanism in the ADMM-based MABO framework.
Major contributions of this work are as follows:
  • We propose a modified DMPC scheme and demonstrate its ability to capture the true multi-agent MDP, even when the local MPC schemes, formulated via the dual decomposition method, are based on imperfect models of the real multi-agent system. We provide formal proofs to establish that the modified DMPC scheme serves as a valid approximator for the local optimal policies.
  • We introduce a parametric variant of the modified DMPC scheme and develop an ADMM-based MABO algorithm to enable coordinated learning of the parameterized local MPC schemes.
  • We conduct a rigorous convergence analysis of the proposed coordination-based distributed learning algorithm, supported by formal proofs of key properties of the acquisition function employed in the ADMM-based MABO method.
@@
This paper is organized as follows. In Section II, a distributed MPC scheme based on a dynamic dual-decomposition method is described. The fusion of DMPC and multi-agent MDPs is detailed in Section III. In Section IV, we provide the proposed coordinated BO algorithm for learning the DMPC schemes. Numerical examples are then given in Section V to show the accuracy, efficacy, and performance of the proposed MABO-DMPC. Concluding remarks are finally given in Section VI.
II Preliminaries and Problem Setup
@@
In this section, we first provide the formulation for the centralized control problem and then describe the use of dual decomposition to break the centralized problem into local (distributed) optimization problems. We consider a distributed networked optimal control system comprising a set of $\mathcal{M}$ interconnected dynamical subsystems, each denoted as =$\mathcal{M}=\{\Sigma_{1},\dots,\Sigma_{|\mathcal{M}|}\}$, where $m:=|\mathcal{M}|$ represents the cardinality of the set $\mathcal{M}$. For each subsystem $\Sigma_{i}$, we define subsystem $\Sigma_{j}$ as a neighbor of $\Sigma_{i}$ if the two subsystems are subject to coupled constraints and/or a coupled cost function. We also consider a Distributed Control System (DCS) defined by a graph $\mathcal{G}(\mathcal{M},\mathcal{E}_{c})$, consisting of a set of nodes $\mathcal{M}$ corresponding to the subsystems and edges $\mathcal{E}_{c}$ representing the interconnections of the subsystems through a coupling cost function. The set $\mathcal{M}_{i}:=\{\Sigma_{j}\,|\,(i,j)\in\mathcal{E}_{c}\text{ or }(j,i)\in\mathcal{E}_{c},\,i\neq j\}$ then represents the set of subsystems $\Sigma_{j}$ that are interconnected with $\Sigma_{i}$.
II-A Centralized Control Problem
@@
Let the states and inputs of the agent $i$ be denoted by $\boldsymbol{\mathrm{x}}_{i}$ and $\boldsymbol{\mathrm{u}}_{i}$, respectively. We consider a deterministic model of each agent as $\boldsymbol{\mathrm{x}}_{i}^{k+1}=\boldsymbol{\mathrm{f}}_{i}(\boldsymbol{\mathrm{x}}_{i}^{k},\,\boldsymbol{\mathrm{u}}_{i}^{k})$, where $\boldsymbol{\mathrm{x}}_{i}^{k}\in\mathbb{R}^{n_{x_{i}}},\boldsymbol{\mathrm{u}}_{i}^{k}\in\mathbb{R}^{n_{u_{i}}}$. Let $\boldsymbol{\mathrm{x}}^{k}=\mathrm{col}\left\{\boldsymbol{\mathrm{x}}_{1}^{k},\ldots,\boldsymbol{\mathrm{x}}_{m}^{k}\right\}$ and $\boldsymbol{\mathrm{u}}^{k}=\mathrm{col}\left\{\boldsymbol{\mathrm{u}}_{1}^{k},\cdots,\boldsymbol{\mathrm{u}}_{m}^{k}\right\}$ be the augmented state and control input vectors of multi-agent system, respectively. The corresponding dimensions then are $\boldsymbol{\mathrm{x}}^{k}\in\mathbb{R}^{n_{x}},\boldsymbol{\mathrm{u}}^{k}\in\mathbb{R}^{n_{u}}$, $n_{x}=\sum_{i=1}^{m}n_{x_{i}},n_{u}=\sum_{i=1}^{m}n_{u_{i}}$. A networked control scheme can be based on a centralized optimization problem or a set of local problems that need to be solved at each time instant. The centralized optimization problem is formulated as
$\min_{\hat{\boldsymbol{\mathrm{x}}},\hat{\boldsymbol{\mathrm{u}}}}\sum_{i=1}^{m}\Bigg{\{}T_{i}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{k+N},\hat{\boldsymbol{\mathrm{w}}}_{i}^{k+N}\right)+\sum_{\ell=k}^{k+N-1}l_{i}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{w}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{u}}}_{i}^{\ell}\right)\Bigg{\}}$ (1a)
@@
where $W_{ij}\left(\hat{\boldsymbol{\mathrm{x}}}_{j}^{\ell}\right)_{(j,i)\in\mathcal{E}_{c}}\in\mathbb{R}^{\Sigma_{j}n_{j}}$ denotes the tuple of the state vector of all subsystems that influence subsystem $i$, $N$ is the prediction horizon, and $T_{i}$, $l_{i}$, $h_{i}$, and $g_{i}$ denote the respective terminal cost, stage cost, mixed inequality constraint, and input inequality constraint for agent $i$, respectively.
@@
In this paper, we assume that the subsystems in the multi-agent system are of the same order (state-space dimension) while their dynamics can be different, yielding a heterogeneous multi-agent system.
@@
In this paper, we consider multi-agent systems with coupled subsystems such that each agent $\Sigma_{i}$ is affected by all other agents $\mathcal{M}_{i}$. According to Assumption 1, we then have that $\hat{\boldsymbol{\mathrm{w}}}_{i}^{\ell}\in\mathbb{R}^{(m-1)n_{x_{i}}}$ in (1d).
Solving (1) gives a sequence of optimal input predictions and corresponding state predictions as
@@
II-B Dynamic Dual Decomposition
Taking into account the coupling constraints (1d) of the centralized optimization problem (1), one can introduce slack variable $\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}$ capturing the effect of other agents on the agent $i$ through $W_{ij}\left(\hat{\boldsymbol{\mathrm{x}}}_{j}^{k+\ell}\right)_{(j,i)\in\mathcal{E}_{c}}$. The centralized optimization scheme can then be reformulated as
@@ II-B Dynamic Dual Decomposition
$+\left(\boldsymbol{\mathrm{\mu}}_{i}^{\ell}\right)^{\top}\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}-\sum_{j=1,j\neq i}^{m}\left(\boldsymbol{\mathrm{\mu}}_{ji}^{\ell}\right)^{\top}W_{ji}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell}\right),$
and $\boldsymbol{\mathrm{\mu}}_{i}^{\ell}=\left(\boldsymbol{\mathrm{\mu}}_{ij}^{\ell}\right)_{(j,i)\in\mathcal{E}_{c}}\in\mathbb{R}^{\Sigma_{j}n_{j}}$. The local MPC problem is then formulated as
@@
$\qquad=\operatorname*{arg\,min}_{\hat{\boldsymbol{\mathrm{x}}}_{i},\hat{\boldsymbol{\mathrm{u}}}_{i},\bar{\boldsymbol{\mathrm{w}}}_{i}}~{}~{}\Bigg{\{}T_{i}^{\boldsymbol{\mathrm{\mu}}^{(I)}}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{k+N},\bar{\boldsymbol{\mathrm{w}}}_{i}^{k+N}\right)$ (5a)
$\qquad\qquad\qquad\qquad\qquad+\sum_{\ell=k}^{k+N-1}L_{i}^{\boldsymbol{\mathrm{\mu}}^{(I)}}\left(\hat{\boldsymbol{\mathrm{x}}}_{i}^{\ell},\hat{\boldsymbol{\mathrm{u}}}_{i}^{\ell},\bar{\boldsymbol{\mathrm{w}}}_{i}^{\ell}\right)\Bigg{\}}$
@@
where $\beta_{i}>0$ is the step size. The stopping criteria are then met ($\boldsymbol{\mathrm{\mu}}_{i}^{I}\rightarrow\boldsymbol{\mathrm{\mu}}_{i}^{\star}$) if either the number of iterations exceeds $I\geq I_{\text{max}}$ or one of the following conditions is satisfied:
@@
for some positive threshold $\epsilon_{1},\epsilon_{2}$. It is worth noting that either one of the two conditions above can be used as a stopping criterion if local MPC models can accurately capture the real subsystems. Consequently, we denote by $\hat{\boldsymbol{\mathrm{\mu}}}_{i}^{\star}$ the local multipliers ($\boldsymbol{\mathrm{\mu}}_{i}^{\star}$ when converged) when coping with imperfect MPC models.